# SWE-RL: Advancing LLM Reasoning via Reinforcement Learning on Open Software Evolution

### v1

@@ 3.2 Main results
Section 3.2 presents the pass@1 results on SWE-bench Verified for models that utilize open-source scaffolds, with models categorized by their size. From the table, we observe that Llama3-SWE-RL-70B achieves state-of-the-art results among small and medium-sized language models (<100B) by resolving 41.0% of the issues. Additionally, all other open-source baselines we compare, such as Lingma-SWE-GPT Ma et al. (2024), SWE-Gym Pan et al. (2024), and SWE-Fixer Xie et al. (2025), include distilled outputs from GPT-4o or Claude-3.5-Sonnet in the training data. In contrast, Llama3-SWE-RL is trained solely with publicly available data through our reinforcement learning technique SWE-RL, without relying on any proprietary LLMs in the pipeline. Llama3-SWE-RL also sets a new record for Llama-based methods on SWE-bench.
{booktabs}
colspec=llrl, Model Scaffold SWE-bench Verified Reference
[…]
@@
Claude-3.5-Sonnet OpenHands 53.0 Wang et al. (2024)
[…]
[c=4]halign=c,font= Model size $\leq$ 100B
@@
SWE-Fixer-72B SWE-Fixer 30.2 Xie et al. (2025)
Llama3-Midtrain-8B (beta)2 Agentless Mini 31.0 Section 8
@@
[…]
[c=4]l 1Open-source Mixture-of-Experts model with 671B total and 37B active parameters
2A beta version developed via extensive midtraining on raw PR data. See Section 8.
@@ 3.3 Baseline comparison
{booktabs}
colspec=llrr, Model Setting Correct format Repair performance (oracle)
@@
To understand how much SWE-RL improves LLMs in solving sofware issues, we compare Llama3-SWE-RL with the corresponding Llama-3 and SFT baseline in Section 3.3, using Agentless Mini as the underlying scaffold. In this experiment, we also evaluate on SWE-bench Verified but focus on the models’ repair ability. To achieve this, we provide oracle files in the context and let the model generate a single repair edit using greedy decoding, without incorporating additional pipeline steps such as localization and test generation. The table reveals that the base Llama-3.3 model struggles to produce correctly formatted code edits, even when using a 20-sample majority voting approach, where outputs with incorrect formats are pre-filtered. With SFT, most code edits generated by the language model are correctly formatted, and the repair performance shows significant improvement. However, Llama3-SWE-RL-70B demonstrates even greater enhancement in repair capabilities, although its format accuracy is slightly lower than that of the SFT version. This indicates that SWE-RL aids the LLM in better reasoning about issue solving and code editing.
@@ 3.4 Scaling analysis with more samples
Figure 5: Scaling analysis with more repair samples and more reproduction tests. The figure on the left illustrates the resolve rate on SWE-bench Verified in relation to the number of repair samples, while maintaining a constant 30 test samples. Conversely, the figure on the right depicts the resolve rate as it varies with the number of reproduction test samples, with a fixed count of 500 repair samples.
Figure 5 shows that increasing both the number of repair samples and test samples enhances performance on SWE-bench. Notably, for repair samples, there is a significant score increase from 33.6 to 40.0 when the sample size is expanded from 20 to 160. However, beyond 160 samples, the improvement trend begins to plateau, with scores only rising slightly from 40.0 to 41.0 as the sample size increases to 320 and 500. Although the impact of adding more reproduction test samples is less obvious, there is still a gradual score improvement from 38.8 to 41.0 as the number of test samples increases up to 20. There is no difference between 20 and 30 test samples, suggesting a performance saturation point has been reached.
@@ 3.5 Generalizability of RL
[…]
Llama3-SWE-RL is only trained with SWE-RL on issue-solving data. This raises a natural question whether such domain-specific training harms the performance on other tasks. To address this, we conduct an experiment in Section 3.5, evaluating the LLMs on five out-of-domain benchmarks, i.e., HumanEval+ Chen et al. (2021); Liu et al. (2023) for function-level code generation, BigCodeBench Zhuo et al. (2024) for practical code generation with library use, CRUXEval Gu et al. (2024) for code execution reasoning, MATH Hendrycks et al. (2021c) for mathematical reasoning, and MMLU Hendrycks et al. (2021b) for general language understanding. We also include the SFT baseline, which is finetuned on the same Llama-3.3-70B-Instruct model using issue-solving data, combined with general coding and dialog data.
@@ 3.6 Reward ablation
[…]
{booktabs}
colspec=lrr, Reward type Correct format Repair (oracle)
@@
As shown in Section 3.6, while the discrete and continuous reward functions lead to similar format accuracy, the continuous reward is more effective in enhancing the repair performance. From the training dynamics, we can see that discrete rewards grow slower than continuous rewards. Additionally, the average discrete reward remains approximately zero upon the completion of training, meaning it struggles to obtain patches exactly matching the oracles. This is because real-world patches are highly diverse and often cannot be easily matched. The continuous reward function better captures partial correctness and incremental improvements, allowing the model to learn more nuanced and effective repair strategies.
@@ 4.1 Language models for software engineering
Large language models (LLMs), trained with billions to trillions of code tokens, have demonstrated outstanding performance in a wide range of coding tasks, including code generation Chen et al. (2021); Li et al. (2023); Rozière et al. (2023); Guo et al. (2024); Wei et al. (2024); Lozhkov et al. (2024); DeepSeek-AI et al. (2024), code optimization Cummins et al. (2024); Liu et al. (2024a), program repair Xia and Zhang (2022); Xia et al. (2023b); Jiang et al. (2023); Wei et al. (2023), and software testing Xia et al. (2023a); Deng et al. (2023); Yuan et al. (2023); Schäfer et al. (2023); Lemieux et al. (2023). Initially, researchers primarily focused on single-shot code generation tasks, such as function-level Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021a); Li et al. (2022); Jain et al. (2024b), class-level Du et al. (2023), and repository-level code completion Ding et al. (2023); Liu et al. (2024b); Zhang et al. (2023). However, with the rapid development of LLMs, the performance on many popular single-shot code generation benchmarks like HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and EvalPlus Liu et al. (2023) has become saturated. Since the development of SWE-bench Jimenez et al. (2023), which requires solving real-world GitHub issues, researchers start to work on improving LLMs’ real-world issue-solving capability and have designed various scaffolds for SWE-bench. Two general types are (1) agentic scaffolds Yang et al. (2024b); Wang et al. (2024); Gauthier (2024), where an LLM drives the decision-making process based on its past actions and observations through tool-based interaction with the environment; and (2) pipeline-based scaffolds Xia et al. (2024); Örwall (2024); Zhang et al. (2024), where an LLM goes through human-defined stages to solve a given issue. Generally, agentic methods are more general but require strong instruction-following and capable LLMs to drive the autonomous process, and can be computationally intensive due to multi-round interactions. In contrast, pipeline-based approaches are more specialized but efficient, with a focus on LLMs’ pure code editing capability. Therefore, we designed our minimalist pipeline-based scaffold, Agentless Mini, to focus on the enhancements of Llama3-SWE-RL’s core code editing capabiltiy.
@@ 5 Conclusion
Limitations. Despite the promising results, our approach has several limitations. First, our reward implementation compares the sequence similarity between the predicted and oracle patch rather than their semantic equivalence. This may prevent the policy LLM from exploring alternative, functional equivalent solutions. Additionally, in Agentless Mini, the localization process is simplified to mapping repository structures to file paths, which lacks comprehensive context. Moreover, as a pipeline-based approach, Agentless Mini divides all steps into distinct inference stages. This “external structure” prevents the model from learning through interaction feedback and hinders its ability to consider the entire problem holistically. Furthermore, our approach requires a substantial sampling budget to achieve optimal results, which may be impractical for projects with high execution costs.
Future work. In the future, we plan to address these limitations by integrating agentic reinforcement learning into our framework. This will enable the model to independently learn localization and repair strategies without relying on external structures. Additionally, we will incorporate execution to allow the model to interact directly with the repository environment. We will also focus on improving the sample efficiency during inference, with the goal of achieving equal or better performance with fewer samples. Ultimately, we aim to enhance the practicality of SWE-RL, paving the way for more powerful, reliable, and fully open-source solutions for active GitHub issue resolution.

### v2

@@ 3.2 Main results
Section˜3.2 presents the pass@1 results on SWE-bench Verified for models that utilize open-source scaffolds, with models categorized by their size. From the table, we observe that Llama3-SWE-RL-70B achieves state-of-the-art results among small and medium-sized language models (<100B) by resolving 41.0% of the issues. Additionally, all other open-source baselines we compare, such as Lingma-SWE-GPT Ma et al. (2024), SWE-Gym Pan et al. (2024), and SWE-Fixer Xie et al. (2025), include distilled outputs from GPT-4o or Claude-3.5-Sonnet in the training data. In contrast, Llama3-SWE-RL is trained solely with publicly available data through our reinforcement learning technique SWE-RL, without relying on any proprietary LLMs in the pipeline. Llama3-SWE-RL also sets a new record for Llama-based methods on SWE-bench. Additionally, we show that our RL-ed model is significantly better than the SFT baseline in this end-to-end setup, indicating that SWE-RL enhances the general reasoning ability of LLMs to solve issues.
Model Scaffold SWE-bench Verified Reference
Model closed-source or size $\gg$ 100B
@@
Claude-3.5-Sonnet OpenHands 53.0 Wang et al. (2024b)
Model size $\leq$ 100B
@@
SWE-Fixer-72B SWE-Fixer 32.8 Xie et al. (2025)
Llama3-SWE-SFT-70B Agentless Mini 36.2 This paper
@@
1Open-source Mixture-of-Experts model with 671B total and 37B active parameters
@@ 3.3 Baseline comparison
Model Setting Correct format Repair performance (oracle)
@@
To understand how much SWE-RL improves LLMs in solving sofware issues, we compare Llama3-SWE-RL with the corresponding Llama-3 and SFT baseline in Table˜2, using Agentless Mini as the underlying scaffold. In this experiment, we also evaluate on SWE-bench Verified but focus on the models’ repair ability. To achieve this, we provide oracle files in the context and let the model generate a single repair edit using greedy decoding, without incorporating additional pipeline steps such as localization and test generation. The table reveals that the base Llama-3.3 model struggles to produce correctly formatted code edits, even when using a 20-sample majority voting approach, where outputs with incorrect formats are pre-filtered. With SFT, most code edits generated by the language model are correctly formatted, and the repair performance shows significant improvement. However, Llama3-SWE-RL-70B demonstrates even greater enhancement in repair capabilities, although its format accuracy is slightly lower than that of the SFT version. This indicates that SWE-RL aids the LLM in better reasoning about issue solving and code editing.
@@ 3.4 Scaling analysis with more samples
Figure 4: Scaling analysis with more repair samples and more reproduction tests. The figure on the left illustrates the resolve rate on SWE-bench Verified in relation to the number of repair samples, while maintaining a constant 30 test samples. Conversely, the figure on the right depicts the resolution rate as it varies with the number of reproduction test samples, with a fixed 500 repair samples.
Figure˜4 shows that increasing both the number of repair samples and test samples enhances performance on SWE-bench. Notably, for repair samples, there is a significant score increase from 33.6 to 40.0 when the sample size is expanded from 20 to 160. However, beyond 160 samples, the improvement trend begins to plateau, with scores only rising slightly from 40.0 to 41.0 as the sample size increases to 320 and 500. Although the impact of adding more reproduction test samples is less obvious, there is still a gradual score improvement from 38.8 to 41.0 as the number of test samples increases up to 20. There is no difference between 20 and 30 test samples, suggesting a performance saturation point has been reached.
@@ 3.5 Generalizability of RL
Llama3-SWE-RL is only trained with SWE-RL on issue-solving data. This raises a natural question whether such domain-specific training harms the performance on other tasks. To address this, we conduct an experiment in Table˜3, evaluating the LLMs on five out-of-domain benchmarks, i.e., HumanEval+ Chen et al. (2021); Liu et al. (2023) for function-level code generation, BigCodeBench Zhuo et al. (2024) for practical code generation with library use, CRUXEval Gu et al. (2024) for code execution reasoning, MATH Hendrycks et al. (2021c) for mathematical reasoning, and MMLU Hendrycks et al. (2021b) for general language understanding. We also include the SFT baseline, which is finetuned on the same Llama-3.3-70B-Instruct model using issue-solving data, combined with general coding and dialog data.
@@
Statistical significance analysis.
While small absolute gains (1-2 percentage points) on HumanEval or CRUXEval are, by themselves, unlikely to be statistically significant, we evaluate across multiple benchmarks and aggregate evidence, observing improvements that consistently favor RL. With Eval Arena Wang et al. (2024a), we indicate that improvements of >0.8>0.8 percentage points on MMLU, 3 points on CRUXEval, and >3>3 points on the full MATH dataset are significant on their own; taken together, these results achieve significance at the $0.05$ level.
@@ 3.6 Reward ablation
According to Equation˜1, the reward design of SWE-RL allows different instantiations of the $\mathit{compare}$ function. Throughout the paper, we adopt the sequence similarity between the predicted and the oracle patch, which is a continuous value from 0 to 1. We denote this type of reward as continuous. It is natural to compare this continuous reward with a discrete reward, where the $\mathit{compare}$ function outputs 1 if the predicted and oracle patches exactly match each other, and 0 otherwise. We trained a variant of Llama3-SWE-RL with the discrete reward function, using the same training setup as in the continuous reward case.
Reward type Correct format Repair (oracle)
@@
As shown in Figure˜5, while the discrete and continuous reward functions lead to similar format accuracy, the continuous reward is more effective in enhancing the repair performance. From the training dynamics, we can see that discrete rewards grow slower than continuous rewards. Additionally, the average discrete reward remains approximately zero upon the completion of training, meaning it struggles to obtain patches exactly matching the oracles. This is because real-world patches are highly diverse and often cannot be easily matched. The continuous reward function better captures partial correctness and incremental improvements, allowing the model to learn more nuanced and effective repair strategies.
@@ 4.1 Language models for software engineering
Large language models (LLMs), trained with billions to trillions of code tokens, have demonstrated outstanding performance in a wide range of coding tasks, including code generation Chen et al. (2021); Li et al. (2023); Rozière et al. (2023); Guo et al. (2024); Wei et al. (2024); Lozhkov et al. (2024); DeepSeek-AI et al. (2024), code optimization Cummins et al. (2024); Liu et al. (2024a), program repair Xia and Zhang (2022); Xia et al. (2023b); Jiang et al. (2023); Wei et al. (2023), and software testing Xia et al. (2023a); Deng et al. (2023); Yuan et al. (2023); Schäfer et al. (2023); Lemieux et al. (2023). Initially, researchers primarily focused on single-shot code generation tasks, such as function-level Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021a); Li et al. (2022); Jain et al. (2024b), class-level Du et al. (2023), and repository-level code completion Ding et al. (2023); Liu et al. (2024b); Zhang et al. (2023). However, with the rapid development of LLMs, the performance on many popular single-shot code generation benchmarks like HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and EvalPlus Liu et al. (2023) has become saturated. Since the development of SWE-bench Jimenez et al. (2023), which requires solving real-world GitHub issues, researchers start to work on improving LLMs’ real-world issue-solving capability and have designed various scaffolds for SWE-bench. Two general types are (1) agentic scaffolds Yang et al. (2024b); Wang et al. (2024b); Gauthier (2024), where an LLM drives the decision-making process based on its past actions and observations through tool-based interaction with the environment; and (2) pipeline-based scaffolds Xia et al. (2024); Örwall (2024); Zhang et al. (2024), where an LLM goes through human-defined stages to solve a given issue. Generally, agentic methods are more general but require strong instruction-following and capable LLMs to drive the autonomous process, and can be computationally intensive due to multi-round interactions. In contrast, pipeline-based approaches are more specialized but efficient, with a focus on LLMs’ pure code editing capability. Therefore, we designed our minimalist pipeline-based scaffold, Agentless Mini, to focus on the enhancements of Llama3-SWE-RL’s core code editing capabiltiy.
@@ 5 Conclusion
Limitations. Despite the promising results, our approach has several limitations. First, our reward implementation compares the sequence similarity between the predicted and oracle patch rather than their semantic equivalence. This may prevent the policy LLM from exploring alternative, functional equivalent solutions. Additionally, in Agentless Mini, the localization process is simplified to mapping repository structures to file paths, which lacks comprehensive context. Moreover, as a pipeline-based approach, Agentless Mini divides all steps into distinct inference stages. This “external structure” prevents the model from learning through interaction feedback and hinders its ability to consider the entire problem holistically.